### v1

# Dynamic Conditional SKEPTIC

Abstract
We introduce the Dynamic Conditional SKEPTIC (DCS), a semiparametric approach for efficiently and robustly estimating time-varying correlations in multivariate models. We exploit nonparametric rank-based statistics, namely Spearman’s rho and Kendall’s tau, to estimate the unknown correlation matrix and discuss the stationarity, $\beta$- and $\rho$- mixing conditions of the model. We illustrate the methodology by estimating the time-varying conditional correlation matrix of the stocks included in the S&P100 and S&P500 during the period from 02/01/2013 to 23/01/2025. The results show that DCS improves diagnostic checks compared to the classical Dynamic Conditional Correlation (DCC) models, providing uncorrelated and normally distributed residuals. A risk management application shows that global minimum variance portfolios estimated using the DCS model exhibit lower turnover than those based on the DCC and DCC-NL models, while also achieving higher Sharpe ratios for portfolios constructed from S&P 100 constituents.
Keywords: Dynamic Conditional Correlation; Latent Gaussian Process; Portfolio Construction; Robust Statistics.
1 Introduction
Multivariate GARCH models analyze the dependence structure of a vector of asset returns through the dynamic estimation of its covariance matrix. When the number of assets is large, these models face computational issues in the estimation procedure that returns poor performance due to the curse of dimensionality. To this end, the recent works of engle2019large and de2022large robustify the Dynamic Conditional Correlation (DCC) (engle2002dynamic), one of the most commonly used in the class of multivariate GARCH models, for estimating time-varying correlations in financial assets, including a nonlinear shrinkage (ledoit2015spectrum) that improves the estimation of the unconditional correlation matrix and relying on the composite likelihood method (pakel2021fitting) that provides computationally feasible estimation of the DCC model in large-dimensional settings.
However, a limitation of the DCC model remains the assumption that asset returns follow a normal distribution. This assumption does not adequately capture the empirical stylized facts (cont2001empirical) which are commonly observed in asset return distributions (bai2003kurtosis), particularly the skewness and leptokurtosis. To formalize the concept, let $\boldsymbol{X}_{t}=(X_{1,t},\ldots,X_{p,t})^{\prime}$ represent the $p$-dimensional random vector of asset returns, the main assumption is that $\boldsymbol{X}_{t}\sim\operatorname{N}(0,H_{t})$, where $H_{t}$ is the covariance matrix. The difference between the normal distribution and the empirical distribution of $\boldsymbol{X}_{t}$ can lead to misleading estimation, resulting in a wrong model fit.
To strengthen the credibility of our thesis, we collect data on the stock returns that compose the S&P 500 index, from 02/01/2013 to 23/01/2025, comprising a total of 429 companies. We perform the Kolmogorov-Smirnov test to compare the empirical distribution of returns against both the normal and Student’s $t$ distributions. For the Student’s $t$ distribution, we estimate the degrees of freedom by fitting a GARCH(1,1) model to the returns, assuming Student’s $t$-distributed innovations.
Table 1: Kolmogorov-Smirnov test results for the normal and Student’s $t$ distributions. In the Table, (N) denotes the normal distribution, while ($t$) represents the Student’s $t$ distribution. The values in the Table are the number of stocks that reject the null hypothesis at 1% and 5% levels.
Significance Level Kolmogorov-Smirnov (N) Kolmogorov-Smirnov ($t$)
1% 429 429
5% 429 429
The test results presented in Table 1 indicate that none of the returns follow either a normal or a Student’s $t$ distribution, highlighting the need for specifying an alternative, more suitable distribution for modeling asset returns.
We introduce the Dynamic Conditional SKEPTIC (DCS) model that relaxes the normality assumption by allowing the marginal distributions to exhibit skewness and heavy tails while employing a Gaussian copula framework to model dependence.
In particular, we consider $\boldsymbol{X}_{t}\sim\operatorname{NPN}(0,H_{t},f)$ where $\operatorname{NPN}$ is the nonparanormal distribution (liu2009nonparanormal) and the set of functions $\{f_{j}\}_{j=1}^{p}$ transform the the random vector $\boldsymbol{X}_{t}=\left(X_{1,t},\ldots,X_{p,t}\right)^{\prime}$, into the random vector $f(\boldsymbol{X}_{t})=\left(f_{1}\left(X_{t,1}\right),\ldots,f_{p}\left(X_{p,t}\right)\right)^{\prime}$, which follows a multivariate Gaussian distribution. The nonparanormal distribution depends on the transformation functions $\{f_{j}\}$ and the covariance matrix $H_{t}$, both estimated from the data.
The nonparanormal approach is particularly appealing when the dimensionality is large, thanks to its robustness against data contamination, driven by extreme market events, that deviate the stock returns from the normal distribution. Robustness is achieved applying monotonic transformations to each marginal, which preserve the original ranking of the variables while aligning the marginal distributions with Gaussian distributions. This allows the nonparanormal model to exploit normal-based dependence structures without requiring normality in the observed data.
When data are truly normally distributed, Spearman’s rho and Kendall’s tau statistics may be used to determine the unknown correlation matrix exploiting their relation with the Pearson correlation coefficient (kruskal1958ordinal). This approach is suggested in the SKEPTIC [^1] Spearman’s rho and Kendall’s tau estimates preempt to infer correlation. estimator (liu2012high), a nonparametric and robust estimator of the correlation matrix that is less sensitive to outliers and does not strictly rely on the normality assumption as the Pearson correlation coefficient does.
We provide a theoretical background to the model, deriving the stationarity conditions and showing the $\beta$- and $\rho$- mixing conditions of the process. We show that our methodology achieves a convergence rate of $O_{P}\left(\sqrt{\log(Tp)/T}\right)$ for the estimation of the unknown correlation matrix.
In the empirical analysis, we estimate the time-varying correlation matrix of the stock returns of companies that compose the S&P 100 and S&P 500 indices and observe that the DCS model leads to a superior fit compared to the DCC and DCC-NL models, enhancing information criteria and diagnostic checks. Additionally, we provide numerical simulations, and we find that the estimates of process parameters are reliable in the case of the DCS model. We also conduct a risk management application, and we construct minimum variance portfolios using the classical Markowitz approach. Notably, the portfolios estimated with the DCS model exhibit low turnover and favorable Sharpe ratios, highlighting both practical benefits and a more robust representation of the underlying dependence structure.
This paper adds to two main research fields: (i) the modeling of temporal dependence using a copula function (beare2010copulas, patton2012review, patton2013copula and fan2014copulas), and (ii) the nonlinear modeling of multivariate time series with a copula-based function, (chen2006estimation, lee2009copula, remillard2012copula, fan2023estimation and cordoni2024consistent). In particular, fan2023estimation consider a semiparametric Gaussian copula process for estimating the transition matrix within a VAR model, whereas cordoni2024consistent use it for causal inference in a high-dimensional time series context.
The outline of the work is the following. In Section 2, we recall the background, whereas the model is presented in Section 3. We provide a theoretical discussion of the model in Section 4 and the estimation procedure in Section 5. We conduct a simulation study in Section 6, whereas in Section 7, we propose an empirical analysis based on Markowitz’s global minimum variance portfolios. Section 8 concludes.
2 Background
2.1 Notation
We assume that $A=\left[A_{ij}\right]\in\mathbb{R}^{p\times p}$ and $m=\left(m_{1},\ldots,m_{p}\right)^{\prime}\in\mathbb{R}^{p}$. For $1\leq q<\infty$, the matrix $\ell_{q}$-operator norm is defined as: $\|A\|_{q}=\sup_{m\neq 0}\frac{\|Am\|_{q}}{\|m\|_{q}}$. For $q=1$ and $q=\infty$, the norms are: $\|A\|_{1}=\max_{1\leq j\leq p}\sum_{i=1}^{p}|A_{ij}|,$ and $\|A\|_{\infty}=\max_{1\leq i\leq p}\sum_{j=1}^{p}|A_{ij}|$. The matrix $\ell_{2}$-operator norm, also known as the spectral norm, is defined as the largest singular value of the matrix $A$. We also define $\|A\|_{\max}=\max_{ij}|A_{ij}|$, representing the maximum absolute value of any element in the matrix. Moreover, $A\geq 0$ means that all elements of $A$ are nonnegative, whereas $A>0$ implies that all elements are strictly positive. The operator $\operatorname{Vech}(A)$ denotes the $p(p+1)/2:=p^{*}$-dimensional column vector formed by stacking the upper triangular elements of $A$ in a column-wise manner, excluding redundancies. We define $\operatorname{diag}(A)=\left[a_{ij}\mathbf{1}(i=j)\right]_{1\leq i,j\leq p}$, and $\operatorname{Vecd}(A)=\left[a_{ii}\right]_{1\leq i\leq p}$, the vector consisting of the diagonal elements of $A$ in $\mathbb{R}^{p}$. We indicate $|A|$ the determinant of the matrix A. In conclusion, $\rho(A)$ represents the spectral radius of $A$, defined as the largest modulus of its eigenvalues. If $A$ is positive semidefinite, then $\rho(A)=\|A\|_{2}$.
2.2 Dynamic Conditional Corelation
Let $\boldsymbol{X}_{t}$ denote a stochastic process, and $\mathcal{F}_{t-1}$ the information set at time t1t-1, we assume that:
    $\mathbb{E}_{t-1}\left(\boldsymbol{X}_{t}\mid\mathcal{F}_{t-1}\right)\equiv\mu_{t}=0$,
    $\mathbb{E}_{t-1}\left(\boldsymbol{X}_{t}\boldsymbol{X}_{t}^{\prime}\mid\mathcal{F}_{t-1}\right)\equiv H_{t}$.
For our purposes, the stochastic process $\boldsymbol{X}_{t}$ represents asset returns, we assume that the asset returns dynamic is:
$\mathbf{X}_{t}=H_{t}^{\frac{1}{2}}\varepsilon_{t}\text{ with }\varepsilon_{t}\sim\mathrm{N}\left(0,I_{p}\right).$ (1)
Consider two generic random variables, $X_{i,t}$ and $X_{j,t}$, their correlation is:
$\rho_{ij,t}=\frac{\mathbb{E}_{t-1}({X}_{i,t}{X}_{j,t})}{\sqrt{\mathbb{E}_{t-1}({X}^{2}_{i,t})\mathbb{E}_{t-1}({X}^{2}_{j,t}})}.$ (2)
We decompose the covariance matrix $H_{t}$ to extrapolate $R_{t}$, the correlation of $\boldsymbol{X}_{t}$:
$H_{t}=D_{t}R_{t}D_{t},\quad\text{ where: }D_{t}^{2}=\text{diag}(H_{t}).$ (3)
To model $R_{t}$, engle2002dynamic assumes that $\boldsymbol{X}_{t}\sim N(0,H_{t})$, where the residuals $\varepsilon_{t}=D_{t}^{-1}\boldsymbol{X}_{t}$ are such that $\mathbb{E}\left(\varepsilon_{t}\varepsilon_{t}^{\prime}\mid\mathcal{F}_{t-1}\right)=R_{t}$ is the unconditional correlation matrix of the returns at time $t$, modeled according to the following linear transformation:
$R_{t}=\operatorname{diag}\left(q_{11,t}^{-\frac{1}{2}}\ldots q_{pp,t}^{-\frac{1}{2}}\right)Q_{t}\operatorname{diag}\left(q_{11,t}^{-\frac{1}{2}}\ldots q_{pp,t}^{-\frac{1}{2}}\right),$ (4)
where $\bar{q}_{ij}\simeq\bar{\rho}_{ij}$, and $\rho_{ij}$ is the element $i,j$ of the correlation matrix of returns:
$\rho_{ij,t}=\frac{q_{ij,t}}{\sqrt{q_{ii,t}q_{jj,t}}}.$ (5)
The Dynamic Conditional Correlation (DCC) model is an autoregressive process given by:
$Q_{t}=\bar{Q}(1-\alpha-\beta)+\alpha(\varepsilon_{t-1}\varepsilon_{t-1}^{\prime})+\beta Q_{t-1},$ (6)
where $\bar{Q}$ is the unconditional covariance matrix of the residuals, and $\alpha$ and $\beta$ are non-negative parameters satisfying $\alpha+\beta<1$. From the residuals $\varepsilon_{i,t}=D^{-1}_{i,t}x_{i,t}=\frac{x_{i,t}}{\sqrt{h_{ii,t}}}$, the estimation of the volatility of the asset return $h_{ii,t}$ over time for each asset $i$ is obtained by fitting GARCH-type models.
In large dimensions, the estimation of the unconditional correlation matrix $\bar{Q}$ suffers from overfitting due to the curse of the dimensionality problem. engle2019large solve this issue with the DCC-NL model, introducing a non-linear shrinkage of the eigenvalues, $\bar{Q}$, using the QuEST function in ledoit2015spectrum that produces a more stable estimate of the matrix.
We propose an extension of the DCC model by relaxing the assumption of normally distributed returns in favor of the more flexible nonparanormal distribution. To capture the evolving dependence structure among asset returns, we employ nonparametric measures of association, specifically Spearman’s rho and Kendall’s tau.
3 The Dynamic Conditional SKEPTIC Model
3.1 Nonparanormal
Let $\boldsymbol{X}_{t}$} denote the demeaned $p$-dimensional vector of asset returns, so that the conditional covariance matrix is $H_{t}=D_{t}R_{t}D_{t}$, where $D_{t}$ collects the univariate conditional volatilities estimated via GARCH-type processes. We define the standardized residuals ${\varepsilon}_{t}=D_{t}^{-1}\boldsymbol{X}_{t}=(\varepsilon_{1t},\ldots,\varepsilon_{pt})^{\prime},$ such that $\mathbb{E}({\varepsilon}_{t}{\varepsilon}_{t}^{\prime}\mid\mathcal{F}_{t-1})=R_{t}$. To relax the assumption of multivariate normality, we assume that the standardized residuals follow a nonparanormal distribution: ${\varepsilon}_{t}\sim\operatorname{NPN}(0,R_{t},f)$, where ${f}=\{f_{1},\ldots,f_{p}\}$ is a collection of strictly increasing univariate functions. For each margin $j=1,\ldots,p$, we define $\nu_{t,j}=f_{j}(\varepsilon_{t,j}),$ so that the component-wise transformation $\nu_{t}=f\left(\varepsilon_{t}\right)=\left(\nu_{t,1},\ldots,\nu_{t,p}\right)^{\prime}$ follows a multivariate normal distribution with zero mean and correlation matrix $R_{t}$: $\nu_{t}\sim\mathrm{N}\left(0,R_{t}\right).$
Under the nonparanormal distribution, the random vector $\boldsymbol{X}_{t}=D_{t}\nu_{t}$ follows the structure $\boldsymbol{X}_{t}\sim\operatorname{NPN}\left(0,H_{t},f\right)$. The joint probability density function of $\boldsymbol{X_{t}}$ is:
$f_{\boldsymbol{X_{t}}}(x_{t})=\frac{1}{(2\pi)^{p/2}|H_{t}|^{1/2}}\exp\!\left\{-\frac{1}{2}f(x_{t})^{\prime}H_{t}^{-1}f(x_{t})\right\}\prod_{j=1}^{p}|f_{j}^{\prime}(x_{t,j})|,$ (7)
where $f(x_{t})=D_{t}f(\varepsilon_{t})$ and $\varepsilon_{t}=D_{t}^{-1}x_{t}$, so that the same conditional scale matrix $D_{t}$ governs both the standardization and the subsequent transformation. so that $f(x_{t})$ applies the monotone transformations to the standardized components of $\boldsymbol{X}_{t}$. We therefore write $\boldsymbol{X}_{t}\sim\operatorname{NPN}(0,H_{t},f).$
In this setting, the conditional covariance structure $H_{t}$ captures the dynamic evolution of the second moments, while the nonlinear transformations $f_{j}$ account for departures from normality in the marginal distributions. Importantly, the dependence structure encoded in $R_{t}$ is not defined directly on the raw variables $\boldsymbol{X}_{t}$, but rather on their standardized counterparts.
Hence, while $\boldsymbol{X}_{t}$ follows a nonparanormal distribution with full covariance $H_{t}=D_{t}R_{t}D_{t}$, the underlying Gaussian copula that governs its dependence structure acts on the standardized innovations ${\varepsilon}_{t}$. Formally, this implies that
$f({\varepsilon}_{t})\sim N(0,R_{t}),\qquad\boldsymbol{Z}_{t}\equiv f(\boldsymbol{X}_{t})=D_{t}f({\varepsilon}_{t})\sim N(0,H_{t}).$
By Sklar’s theorem, since the $f_{j}$’s are monotone and differentiable, the joint distribution of ${\varepsilon}_{t}$ corresponds to a semiparametric Gaussian copula with correlation $R_{t}$:
$f\left({\varepsilon}_{t}\right)=f\left(\left(\varepsilon_{1t},\ldots,\varepsilon_{pt}\right)^{\prime}\right)=\Phi_{R_{t}}\left(\Phi^{-1}\left(F_{1}\left(\varepsilon_{1t}\right)\right)^{\prime},\ldots,\Phi^{-1}\left(F_{p}\left(\varepsilon_{pt}\right)\right)^{\prime}\right).$
where $\Phi_{R_{t}}$ denotes the multivariate Gaussian cumulative distribution function (CDF) with correlation matrix $R_{t}$.
To estimate the copula function over a time series of length $T$, we use the empirical cumulative distribution function (ECDF) of the standardized residuals:
$\widehat{F}_{j}(w)=\frac{1}{T+1}\sum_{i=1}^{T}\mathbb{I}(\varepsilon_{j,i}\leq w),\qquad\widehat{f}_{j}(\varepsilon_{j,t})=\Phi^{-1}\!\big(\widehat{F}_{j}(\varepsilon_{j,t})\big),$
which yields the nonparametric normal scores $\widehat{{\nu}}_{t}=(\widehat{f}_{1}(\varepsilon_{1t}),\ldots,\widehat{f}_{p}(\varepsilon_{pt}))^{\prime}$.
fan2023estimation and cordoni2024consistent extend the nonparanormal framework to dependent data, proposing vector autoregressive structures for latent Gaussian scores. In our setting, we focus instead on modeling the second-order dynamics through $R_{t}$, leading to the Dynamic Conditional SKEPTIC specification.
3.2 Dynamic Conditional SKEPTIC
Under the nonparanormal assumption, we estimate the time-varying correlation matrix $R_{t}$ using rank-based dependence measures, specifically Spearman’s rho and Kendall’s tau.
Lemma 3.1.
Assume that $\boldsymbol{X}_{t}\sim\operatorname{NPN}(0,H_{t},f)$, where $H_{t}=D_{t}R_{t}D_{t}$. Then the elements of the correlation matrix $R_{t}$ satisfy:
$R_{ij,t}]=2\sin\left(\frac{\pi}{6}\rho^{s}_{ij,t}\right)=\sin\left(\frac{\pi}{2}\tau_{ij,t}\right),$
where $\rho^{s}_{ij,t}$ and $\tau_{ij,t}$ denote Spearman’s rho and Kendall’s tau between the variables $X_{it}$ and $X_{jt}$, respectively.
Motivated by this result, we define two estimators of the dynamic correlation matrix $R_{t}$, based on the observations available up to time $t$. These estimators are denoted by $\widehat{R}^{\rho}_{t}=[\widehat{R}^{\rho}_{ij,t}]$ and $\widehat{R}^{\tau}_{t}=[\widehat{R}^{\tau}_{ij,t}]$, with individual entries defined as:
$\widehat{R}^{\rho}_{ij,t}=\begin{cases}2\sin\left(\frac{\pi}{6}\widehat{\rho}^{s}_{ij,t}\right)&\text{if }i\neq j\\ 1&\text{if }i=j\end{cases},\qquad\widehat{R}^{\tau}_{ij,t}=\begin{cases}\sin\left(\frac{\pi}{2}\widehat{\tau}_{ij,t}\right)&\text{if }i\neq j\\ 1&\text{if }i=j\end{cases}.$
Given a multivariate time series $\boldsymbol{X}_{t}=\left(X_{1t},\ldots,X_{pt}\right)$, the time-varying Kendall’s tau statistic is defined as:
$\widehat{\tau}_{ij,t}:=\frac{2}{T(T-1)}\sum_{1\leq s<s^{\prime}\leq T}\operatorname{sign}(X_{is}-X_{is^{\prime}})\operatorname{sign}(X_{js}-X_{js^{\prime}}),$
while Spearman’s rho is:
$\widehat{\rho}^{s}_{ij,t}:=\frac{\sum_{s=1}^{T}(\zeta_{is}-\bar{\zeta}_{i})(\zeta_{js}-\bar{\zeta}_{j})}{\sqrt{\sum_{s=1}^{T}(\zeta_{is}-\bar{\zeta}_{i})^{2}}\sqrt{\sum_{s=1}^{T}(\zeta_{js}-\bar{\zeta}_{j})^{2}}},$
where $\zeta_{is}$ and $\zeta_{js}$ are the ranks of $X_{is}$ and $X_{js}$ among the observations $(X_{i1},\dots,X_{iT})$ and $(X_{j1},\dots,X_{jT})$, respectively, and $\bar{\zeta}_{i}=\bar{\zeta}_{j}=(T+1)/2$ denotes the average rank. For a generic time $t$, these statistics provide nonparametric estimates of the dependence between the empirical realizations of $X_{it}$ and $X_{jt}$. Let $\widetilde{X}_{it}$ and $\widetilde{X}_{jt}$ be independent copies of $X_{it}$ and $X_{jt}$, and let $\Phi_{i}$ and $\Phi_{j}$ denote their marginal distribution functions. Then, the population versions of Spearman’s rho and Kendall’s tau are defined as: $\rho^{s}_{ij}:=\operatorname{Corr}\left(\Phi_{i}(X_{it}),\Phi_{j}(X_{jt})\right)$, and $\tau_{ij}:=\operatorname{Corr}\left(\operatorname{sign}(X_{it}-\widetilde{X}_{it}),\operatorname{sign}(X_{jt}-\widetilde{X}_{jt})\right).$ Substituting $\widehat{R}^{\rho}_{t}$ or $\widehat{R}^{\tau}_{t}$ to the canonical correlation matrix used in engle2002dynamic, we introduce the Dynamic Conditional SKEPTIC.
Definition 3.2.
The Dynamic Conditional SKEPTIC (DCS) is defined in (3), where $R_{t}$ is estimated according to Lemma 3.1: $\mathcal{R}_{t}=\{{R}^{\rho}_{t}\vee{R}^{\tau}_{t}\}$. Specifically,
$\mathcal{R}_{t}=\operatorname{diag}\left(q_{11,t}^{-\frac{1}{2}},\ldots,q_{pp,t}^{-\frac{1}{2}}\right)\mathcal{Q}_{t}\operatorname{diag}\left(q_{11,t}^{-\frac{1}{2}},\ldots,q_{pp,t}^{-\frac{1}{2}}\right),$ (8)
where the $p\times p$ symmetric positive definite matrix $\mathcal{Q}_{t}=(q_{ij,t})$ is given by:
$\mathcal{Q}_{t}=(1-\alpha-\beta)\bar{\mathcal{Q}}+\alpha{\nu}_{t-1}{\nu}_{t-1}^{\prime}+\beta\mathcal{Q}_{t-1},$ (9)
with $\nu_{it}=f(\varepsilon_{it})$. The matrix $\bar{\mathcal{Q}}$ represents the $p\times p$ unconditional variance matrix of $\nu_{t}$, estimated using SKEPTIC. The parameters $\alpha$ and $\beta$ are non-negative scalars satisfying the constraint $\alpha+\beta<1$.
4 Theoretical Properties
We establish the theoretical foundation of the SKEPTIC estimator for an autoregressive process, demonstrating that this novel approach is a reliable extension to the DCC model. We start recalling the stationarity conditions outlined in the work of fermanian2017stationarity, where the DCC is reformulated in terms of a Markov Chain. In our case, under the assumption of nonparanormal distribution, we define the following model:
$W_{t}=\left(W_{t}^{(1)},W_{t}^{(2)},W_{t}^{(3)},W_{t}^{(4)}\right)^{\prime}\quad\text{where}$
$W_{t}^{(1)}=\left(\operatorname{Vecd}\left(D_{t}\right)\right)^{\prime},$ $W_{t}^{(2)}=\left(\vec{X}_{t}\right)^{\prime},\ \text{where }\vec{X}_{t}=(X_{1t}^{2},\ldots,X_{pt}^{2})^{\prime},$
$W_{t}^{(3)}=\left(\operatorname{Vech}\left(\mathcal{Q}_{t}\right)\right)^{\prime},$ $W_{t}^{(4)}=\left(\operatorname{Vech}\left(\nu_{t}\nu^{\prime}_{t}\right)\right)^{\prime}.$
The dimensions of the four previous random vectors are, respectively, $1\times p,1\times p,1\times p(p+1)/2$, and $1\times p(p+1)/2$. Their sum, dimension of $W_{t}$, is denoted by $d$. Simple block matrix calculations show that there exist random matrices, $\left(T_{t}\right)$, and a vector process, $\left(\kappa_{t}\right)$, such that the dynamics of $W_{t}$ may be rewritten as:
$W_{t}=T_{t}\cdot W_{t-1}+\kappa_{t},$ (10)
for any $t$, where $T:=[T_{ij,t}]_{1\leq i,j\leq 4}$ is a block matrix. [^2] We refer to fermanian2017stationarity, where $\mu=\nu=r=s=1$. To determine the stationarity of the DCS process, we define the data-generating process of $W_{t}$:
$\delta_{t}:=\mathcal{R}^{-\frac{1}{2}}\nu_{t}=\mathcal{R}^{-\frac{1}{2}}D^{-\frac{1}{2}}W_{t}.$ (11)
Note that $\mathbb{E}\left[\delta_{t}\mid\mathcal{F}_{t-1}\right]=0$ and $\mathbb{E}\left[\delta_{t}\delta_{t}^{\prime}\mid\mathcal{F}_{t-1}\right]=I_{p}$ by construction. The definition of these innovations implies that, for every $t,\sigma\left(\delta_{j},j\leq t\right)\subset\sigma\left(\nu_{j},j\leq t\right)\subset\mathcal{F}_{t}$, where $\sigma$ is the sigma algebra.
Proposition 4.1.
(Stationarity of DCC (fermanian2017stationarity)) Given the process $\delta_{t}$ in (11), we state that:
  • (A0)
    $\left(\delta_{t}\right)_{t\in\mathbb{Z}}$ possesses the Markov property with respect to the filtration $\mathcal{F}$. In particular, $\mathbb{E}\left[\delta_{t}\mid\mathcal{F}_{t-1}\right]=E\left[\delta_{t}\mid W_{t-1}\right]$ for every $t$.
  • (A1)
    $\left\|M_{1}\right\|_{\infty}+\left\|N_{1}\right\|_{\infty}<1\quad\text{and}\quad\left\|B\right\|_{\infty}<1.$
    where: $Vecd(D_{t})=V_{0}+M_{1}Vecd(D_{t-1})+N_{1}\vec{W}_{t}$ are the univariate GARCH(1,1) models, and the matrix $B\geq 0$, in other words, all the elements must be non-negative.
  • (A2)
    For some $\mathbb{E}\left[\left\|\delta_{t}\right\|^{2}\right]<\infty$ and $\rho\left(T^{*}\right)<1$, where $T^{*}:=\sup_{\mathbf{w}\in\mathbb{R}^{d}}E\left[\left|T_{t}\right|\mid W_{t-1}=\mathbf{w}\right]$, for any norm $\|\cdot\|$.
  • (A3)
    The law of $\delta_{t}$ given that $W_{t-1}=\mathbf{w}$ is continuous concerning the Lebesgue measure, and its density is denoted by $f_{\delta_{t}}(\cdot\mid\mathbf{w})$, for every $\mathbf{w}\in\mathbb{R}^{d}$ and $t$. The function $\mathbf{w}\mapsto f_{\delta_{t}}(\delta\mid\mathbf{w})$ is continuous for every $\delta\in\mathbb{R}^{p}$ and $t$. There exists an integrable function $H$ s.t. $\sup_{t}\sup_{\mathbf{w}\in\mathbb{R}^{d}}f_{\delta_{t}}(\delta\mid\mathbf{w})\leq H(\delta)$ for every $\delta\in\mathbb{R}^{p}$. Moreover, $\sup_{t}E\left[\left\|\delta_{t}\right\|^{2}\mid W_{t-1}=\mathbf{w}\right]\leq\bar{h}(\|\mathbf{w}\|)$, for some function $\bar{h}$ that satisfies $\lim_{v\rightarrow+\infty}\bar{h}(v)/v^{\gamma}=0$ for every $\gamma>0$.
Under the assumptions (A0, A1, A2, A3), the process $(W_{t},D_{t},\mathcal{R}_{t})$ is stationary.
When the assumptions are satisfied, the correlation matrix estimated with the Dynamic Conditional SKEPTIC in Definition 3.2 is an autoregressive process, which is a stationary Markov chain.
4.1 Mixing Conditions of Semiparametric Gaussian Process
Under the assumption of $\boldsymbol{X}_{t}\sim\operatorname{NPN}(0,H_{t},f)$. For each sequence of random variables, we derive its mixing properties.
Definition 4.2.
(beare2010copulas) The $\beta$-mixing coefficients $\left\{\beta_{k}:k\in\mathbb{N}\right\}$ that correspond to the sequence of random variables $\left\{X_{t}\right\}$ are given by
$\beta_{k}=\frac{1}{2}\sup_{m\in\mathbb{Z}}\sup_{\left\{A_{i}\right\},\left(B_{j}\right\}}\sum_{i=1}^{I}\sum_{j=1}^{J}\left|P\left(A_{i}\cap B_{j}\right)-P\left(A_{i}\right)P\left(B_{j}\right)\right|$
where the second supremum is taken over all finite partitions $\left\{A_{1},\ldots,A_{I}\right\}$ and $\left\{B_{1},\ldots,B_{J}\right\}$ of $\Omega$ such that $A_{i}\in\mathcal{F}_{-\infty}^{m}$ for each $i$ and $B_{j}\in\mathcal{F}_{m+k}^{\infty}$ for each $j$.
Proposition 4.3.
The semiparametric Gaussian copula $C$ is symmetric and continuous with square-integrable density $c$ with $\rho_{C}<1$, then there exists $A<\infty$ and $\gamma>0$ such that $\beta_{k}\leq Ae^{-\gamma k}$ for all $k$.
The proof is in the Appendix A. Moreover, by Proposition 2, and 4 in remillard2012copula, the Gaussian copula process is also a $\rho$-mixing process, where the $\rho$-mixing coefficients are defined as follows.
Definition 4.4.
(beare2010copulas) The $\rho$-mixing coefficients $\left\{\rho_{k}:k\in\mathbb{N}\right\}$ that correspond to the sequence of random variables $\left\{X_{t}\right\}$ are given by
$\rho_{k}=\sup_{m\in\mathbb{Z}}\sup_{f,g}|\operatorname{Corr}(f,g)|,$
where the second supremum is taken over all square-integrable random variables $f$ and $g$ measurable for $\mathcal{F}_{-\infty}^{m}$ and $\mathcal{F}_{m+k}^{\infty}$ respectively, with positive and finite variance, and where $\operatorname{Corr}(f,g)$ denotes the correlation between $f$ and $g$.
The semiparametric Gaussian copula exhibits strong mixing properties, characterizing a process where serial dependence among random variables diminishes over time. We exploit these mixing conditions to derive concentration inequalities for the Dynamic Conditional SKEPTIC. We recall that Spearman’s rho and Kendall’s tau can be expressed in terms of U-statistics. The U-statistic of Kendall’s tau is:
$\tau_{2}\left({X}_{1},\ldots,{X}_{T}\right):=\frac{2}{T(T-1)}\sum_{t<t^{\prime}}\operatorname{sign}\left(X_{t1}-X_{t^{\prime}1}\right)\operatorname{sign}\left(X_{t2}-X_{t^{\prime}2}\right),$
whereas for Spearman’s rho we have:
$\rho=$ $\frac{T-2}{T+1}\cdot\underbrace{\frac{3}{T(T-1)(T-2)}\sum_{t\neq t^{\prime}\neq t^{\prime\prime}}\operatorname{sign}\left(X_{t1}-X_{t^{\prime}1}\right)\operatorname{sign}\left(X_{t2}-X_{t^{\prime\prime}2}\right)}_{\rho_{3}\left(\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{T}\right)}$
$+\frac{3\tau_{2}\left(X_{1},\ldots,{X}_{T}\right)}{T+1},$
where $\tau_{2}(X_{1},\ldots,X_{T})$ is a U-statistic of order two, and $\rho_{3}\left({X}_{1},\ldots,{X}_{T}\right)$ is a U-statistic of order three with an asymmetric kernel function. Under Definition 4.2 and Proposition 4.3, we obtain bounds for our estimators.
Proposition 4.5.
Suppose that the Definition 4.2 and the Proposition 4.3 hold, and assume $(W_{t},D_{t},\mathcal{R}_{t})$ is a stationary process, then we have
$\|\widehat{R}^{\tau}_{t}-R_{t}\|_{\max}=O_{P}(\sqrt{\log(Tp)/T})\quad\text{and}\quad\|\widehat{R}^{\rho}_{t}-R_{t}\|_{\max}=O_{P}(\sqrt{\log(Tp)/T}).$ (12)
This result is crucial, as it demonstrates that, under mixing conditions and stationarity, the estimator of the conditional correlation matrix remains robust even as the dimensionality increases. The proof is in Appendix B.
5 Estimation
The parameters estimation procedure in large dimensions is burdensome due to the difficulty in inverting the matrix $H_{t}$ in the log-likelihood of $z_{t}\equiv f(x_{t})\mid\mathcal{F}_{t-1}\sim\operatorname{N}\left(0,D_{t}\mathcal{R}_{t}D_{t}\right)$:
$L=\sum_{t=1}^{T}\left(\frac{1}{2}\log\left|H_{t}\right|+\frac{1}{2}z_{t}^{\prime}H_{t}^{-1}z_{t}+\sum_{j=1}^{p}\log\left(\left|z_{tj}\right|\right)\right),$
which is proportional to:
$L\propto\frac{1}{2}\sum_{t=1}^{T}\left(\log\left|H_{t}\right|+z_{t}^{\prime}H_{t}^{-1}z_{t}\right).$ (13)
We rewrite the function to extrapolate the correlation term and the volatility terms:
$L$ $=\frac{1}{2}\sum_{t=1}^{T}\left(\log\left|H_{t}\right|+z_{t}^{\prime}H_{t}^{-1}z_{t}\right),$
$=\frac{1}{2}\sum_{t=1}^{T}\left(\log\left|D_{t}\right|^{2}+\log\left|\mathcal{R}_{t}\right|+\nu_{t}^{\prime}\mathcal{R}_{t}\nu_{t}\right)\quad\text{where: }\nu_{t}=f(\varepsilon_{t}),$
$=\frac{1}{2}\sum_{t=1}^{T}\left(\log\left|D_{t}\right|^{2}+z_{t}^{\prime}D^{-1}_{t}D^{-1}_{t}z_{t}-\nu_{t}^{\prime}\nu_{t}+\log\left|\mathcal{R}_{t}\right|+\nu_{t}^{\prime}\mathcal{R}_{t}\nu_{t}\right).$
Denote with $\theta$ the parameters in $D_{t}$ and with $\phi$ the parametrs in $\mathcal{R}_{t}$. The log-likelihood function can be expressed as the sum of a volatility component and a correlation component:
$L(\theta,\phi)=L_{V}(\theta)+L_{C}(\theta,\phi).$
where the volatility term is
$L_{V}(\theta)=\frac{1}{2}\sum_{t=1}^{T}\left(\log\left|D_{t}\right|^{2}+z_{t}^{\prime}D_{t}^{-2}z_{t}\right),$ (14)
and the correlation component is
$L_{C}(\theta,\phi)=\frac{1}{2}\sum_{t=1}^{T}\left(\log\left|\mathcal{R}_{t}\right|+\nu_{t}^{\prime}R_{t}^{-1}\nu_{t}+\nu_{t}^{\prime}\nu_{t}\right).$ (15)
Fitting individual GARCH-type models maximizes the likelihood function for the volatility component. Once the volatility parameters are determined, we need to perform the second step of the estimation procedure, which requires maximizing $\phi$ given $\hat{\theta}$:
$\max_{\phi}\left\{L_{C}(\hat{\theta},\phi)\right\}.$ (16)
The estimation of $\phi=(\alpha,\beta)$ proposed in engle2002dynamic leads to major issues as dimensionality grows. This difficulty aligns with the curse of dimensionality described in engle2019large. Specifically, as the concentration ratio $p/T$ increases, the estimation of $\mathcal{R}_{t}$ suffers from overfitting, and when $p/T$ is equal to 1, the matrix $\mathcal{R}_{t}$ becomes singular and the procedure inapplicable.
We follow pakel2021fitting and use the composite likelihood function that estimates the parameters by summing the quasi-likelihoods of asset subsets, thereby addressing lower-dimensional marginal densities. This method involves determining the conditional variances with a GARCH-type model and then fitting a likelihood based on correlation $2\times 2$. More precisely, once the values of $\hat{\theta}$ that maximize (14) are obtained (fermanian2024model), we estimate the parameter set $\phi$ maximizing:
$L_{C,2T}(\hat{\theta},\phi)=\frac{1}{T}\sum_{t=1}^{T}\sum_{k=1}^{\mathcal{L}}\left[\log\left(\left|\mathcal{R}_{t}^{(k)}\right|\right)+\nu_{t}^{(k)\prime}\left(\mathcal{R}_{t}^{(k)}\right)^{-1}\nu_{t}^{(k)}\right],$ (17)
where $\mathcal{R}_{t}^{(k)}$ is the $2\times 2$ correlation matrix estimated with SKEPTIC with $k$ corresponding to a prespecified pair of indices in $\{1,\ldots,N-1\}$, $\nu_{t}^{(k)}$ is the corresponding $2\times 1$ sub-vector of the standardized residuals $\nu_{t}:=D_{t}^{-1/2}z_{t}$. We choose contiguous pairs $\nu_{t}^{(1)}=\left(\nu_{1t},\nu_{2t}\right)^{\prime},\nu_{t}^{(2)}=$ $\left(\nu_{2t},\nu_{3t}\right)^{\prime},\ldots,\nu_{t}^{(N-1)}=\left(\nu_{(N-1)t},\nu_{Nt}\right)^{\prime}$. By restricting attention to $2\times 2$ matrices, we avoid the need to invert large-dimensional correlation matrices, reducing both computational complexity and the risk of estimation inaccuracies.
5.1 Asymptotic Properties of Composite Likelihood Estimator
Let us recall a bivariate DCS model for $x_{t}\sim\operatorname{NPN}(0,H_{t})$:
$\mathcal{Q}_{t}=(1-\alpha-\beta)\bar{\mathcal{Q}}+\alpha(\nu_{t-1}\nu_{t-1}^{\prime})+\beta\mathcal{Q}_{t-1},$ (18)
where:
$X_{jt}=\left\{x_{1jt},x_{2jt}\right\},\quad\operatorname{Cov}\left(X_{jt}\mid\mathcal{F}_{t-1}\right)=\left(\begin{array}[]{cc}h_{1jt}^{1/2}&0\\ 0&h_{2jt}^{1/2}\end{array}\right)\mathcal{R}_{jt}\left(\begin{array}[]{cc}h_{1jt}^{1/2}&0\\ 0&h_{2jt}^{1/2}\end{array}\right),$
where $\mathrm{Var}\left(x_{ljt}\mid\mathcal{F}_{t-1},\eta_{lt}\right)=h_{ljt}$, $l=1,2$, $\mathrm{Cor}\left(X_{jt}\mid\mathcal{F}_{t-1}\right)=\mathcal{R}_{jt}$, $\mathcal{R}=\left\{R_{jt}^{\rho}\vee R_{jt}^{\tau}\right\}$ and $\theta_{j}=(\eta^{\prime}_{1j},\eta^{\prime}_{2j})^{\prime}$ are estimated with a GARCH-type model. We estimate the parameters: $\phi=(\alpha,\beta)$ and the matrix $\bar{\mathcal{Q}}$ with SKEPTIC in Lemma 3.1. The generic framework for performing the composite likelihood estimation procedure is the following:
$\hat{\theta}=\arg\max_{\phi\in\Phi}\sum_{j=1}^{N}\sum_{t=1}^{T}\ell_{jt}\left(\theta,\hat{\theta}_{j}(\phi)\right)\quad\text{ and }\quad\hat{\lambda}_{j}(\theta)=\arg\max_{\theta_{j}\in\Theta_{j}}\sum_{t=1}^{T}\ell\left(\phi,\theta_{j};X_{jt}\right).$ (19)
pakel2021fitting demonstrate that the composite likelihood estimator yields consistent and asymptotically normal estimates.
6 Simulation Studies
We explore the properties of the Dynamic Conditional SKEPTIC by conducting Monte Carlo simulations. To make the simulations realistic, we first estimate the unconditional covariance matrix of the S&P 500 constituents, denoted by $H_{t}$. We then generate a sample of 1000 simulated returns according to the following dynamics. Let $\boldsymbol{X}_{t}$ denote the vector of returns; we assume
$x_{t}=H^{1/2}_{t}u_{t},$ (20)
where $u_{t}\sim t_{\nu}\left(0,\sigma_{t}\right)$, and the number of degrees of freedom $\nu$ is randomly drawn between 4 and 8. As is well known, empirical returns exhibit non-negligible skewness. To reproduce this feature, we introduce contamination into the simulated data. Let $\delta$ denote the contamination level, we consider three cases, $\delta\in$ {0.01,0.05,0.100.01,0.05,0.10}, corresponding to $1\%,5\%$, and $10\%$ of contaminated observations, respectively. The contamination is introduced by replacing the specified proportion of the data with draws from a $t_{\nu}\left(0,\sigma_{t}\right)$ distribution with $\nu=3$, thereby increasing the number of outliers and inducing asymmetry in the simulated return distribution. We then simulate a DCC model with parameters $\alpha=0.02$ and $\beta=0.97$ as in pakel2021fitting.
For each level of $\delta$, we estimate the parameters of the DCC-based processes and the DCS models $\{\alpha,\beta\}$. We perform the simulations 100 times. We consider four different dimensions: $p\in\{50,97,250,429\}$, where $p=429$, and $p=97$ correspond respectively to the dimension of the S&P500 and S&P100 constituents. In the cases of $p=50$, and $p=250$ we randomly select stocks from the constituents of S&P500 to estimate $H_{t}$. The number of observations we simulate are: $T\in\{500,1000,2000\}$. The results of the simulation are shown in Appendix C. The simulation results highlight that the persistence parameter $\beta$ in the DCS models is lower than the DCC-based specification. In contrast, the $\alpha$ parameters are average higher in the DCS models and closer to the true values than those obtained from the DCC model. The larger $\alpha$ coefficients suggest that the DCS model responds more promptly to short-term fluctuations in dependence. Conversely, the DCC and DCC-NL models, which rely on Gaussian assumptions, tend to smooth out such variations and exhibit a slower adjustment to shocks.
It is worth noting to highlight that as the level of contamination, explained by $\delta$ increases, the standard deviation of the estimates of the DCS model is lower than the DCC-based models, highlighting the ability of the Dynamic Conditional SKEPTIC to mitigate the impact of outlier values.
7 Empirical Analysis
We conduct an empirical analysis comparing the Dynamic Conditional SKEPTIC (DCS) with the DCC model of engle2002dynamic and the DCC-NL of engle2019large [^3] We thank Michael Wolf for kindly providing the codes for the DCC and DCC-NL on the website https://www.econ.uzh.ch/en/people/faculty/wolf/publications.html under the link “Programming Code.” We revisit these codes, aligning them to our purposes.. We download stock price data from Yahoo! Finance of the stocks that belong to the S&P100 and S&P500 indices, covering the period from 02/01/2013 to 23/01/2025. After cleaning the data, the final dataset includes 429 stocks for the S&P500 and 97 stocks for the S&P100, each with 3034 price observations.
7.1 Model Fit and Diagnostic Checks
Using the full set of observations, we fit the DCC, DCC-NL, and DCS models to the S&P100 and S&P500 datasets. Model selection is evaluated according to the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC). For all the models, after having estimated the univariate volatility with a GARCH(1,1), we estimate the parameters $\phi=(\alpha,\beta)$. We remark that the estimate of the matrix ${Q}_{t}$ may not be positive semidefinite. In that case, we project it onto the nearest correlation matrix using the method of higham2002computing.
In the residual analysis, we evaluate the adequacy of model fitting to the data. A well-specified econometric model should effectively filter the observed data into a white noise series, uncorrelated over time. To assess the absence of correlation, we rely on the Portmanteau test (lutkepohl2005new), which examines the null hypothesis that residuals exhibit no autocorrelation up to lag $h$. The test statistic is defined as follows:
$Q_{h}:=T\sum_{i=1}^{h}(T-i)^{-1}\text{tr}(\hat{C}_{i}\hat{C}_{0}^{-1}\hat{C}_{i}\hat{C}_{0}^{-1}),$
where for the DCC and DCC-NL models we have:
$\hat{C}_{i}=\frac{1}{T}\sum_{t=i+1}^{T}\hat{\epsilon}_{t}\hat{\epsilon}_{t-i}^{\prime},$
and for the DCS:
$\hat{C}_{i}=\frac{1}{T}\sum_{t=i+1}^{T}\hat{\nu}_{t}\hat{\nu}_{t-i}^{\prime}.$
In large samples, and for sufficiently large $h$, the test statistic follows an asymptotic Chi-squared distribution:
$Q_{h}\approx\chi^{2}_{(p^{2}(h-k))},$
where $p$ is the dimension of the residuals vector, $h$ is the number of lags considered, and $k$ is the number of estimated parameters. Furthermore, we assess the normality of the residuals using the Jarque-Bera test based on the null hypothesis of normality, where the test statistic is derived from the third and fourth moments of the normal distribution, which asymptotically follows a $\chi^{2}_{2}$ distribution.
Table 2 summarizes the diagnostic checks. For the Portmanteau test, we set the number of lags $h=10$.
S&P100 S&P500
Model AIC BIC JB Port. AIC BIC JB Port.
DCC -29.23 -17.20 $69\times 10^{6\text{ }***}$ 49.03 -28.97 -16.94 $358\times 10^{6\text{ }***}$ 1071.70
DCC-NL -29.23 -17.20 $67\times 10^{6\text{ }***}$ 49.14 -28.97 -16.93 $361\times 10^{6\text{ }***}$ 1084.73
DCS (Tau) -29.32 -17.29 $146.73$ 32.85 -29.06 -17.02 $930.04$ 655.57
DCS (Rho) -29.22 -17.18 $106.92$ 33.43 -29.00 -16.93 $549.08$ 636.29
Table 2: Diagnostic checks of the DCC, DCC-NL, and DCS models fitted to S&P100 and S&P500 dataset. JB stands for Jacque-Bera test statistic values, and Port. stands for Portmanteau test statistic values. Note: $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
We find that the AIC and BIC suggest that our model shows a slight improvement to fit the data, the resulting residuals are uncorrelated for all the models, and in the case of DCS models they satisfy the normal distribution, as highlighted in the results of the Jacque Bera test, where the null hypothesis of normality is not rejected.
7.2 Portfolio Construction
We evaluate the out-of-sample performance of our correlation matrix estimators, which are used to construct the weights of Markowitz portfolios. Each dataset is divided into an in-sample period and an out-of-sample period. The in-sample period includes the first 1.500 observations, from 02/01/2013 to 14/12/2018. The out-of-sample period considers returns from 15/12/2018 to 23/01/2025, consisting of 73 months, where we adopt the convention that 21 consecutive trading days constitute one month. During each month portfolio weights are fixed and no transactions occur. We denote the investment dates by $h=1,\ldots,73$. On each investment date $h$, the covariance matrix is estimated using the most recent $T=1.500$ daily returns, roughly corresponding to six years of historical data.
We estimate the global minimum variance (GMV) portfolio without short-selling constraints. Let ${H}_{t}$ denote the covariance matrix at time $t$. The optimization problem is formulated as follows:
$\min_{w}\,w^{\prime}H_{t}w\quad\text{s.t.}$ $\quad w^{\prime}\mathbf{\iota}=1,$
where $\mathbf{\iota}$ is a vector of ones with dimensions $p\times 1$. The analytical solution is given by:
$w_{t}=\frac{H_{t}^{-1}\mathbf{\iota}}{\mathbf{\iota}^{\prime}H_{t}^{-1}\mathbf{\iota}}.$ (21)
In practice, the unknown $H_{t}$ is replaced with an estimator $\widehat{H}_{t}$, resulting in the feasible portfolio:
$\widehat{w}_{t}=\frac{\widehat{H}_{t}^{-1}\mathbf{\iota}}{\mathbf{\iota}^{\prime}\widehat{H}_{t}^{-1}\mathbf{\iota}}.$ (22)
We evaluate the following five portfolios, constructed using different covariance matrix estimators:
  1. 1.
    DCC: the portfolio is defined in (21), where $\widehat{H}_{t}$ is obtained from a Dynamic Conditional Correlation (DCC) model based on the sample correlation matrix.
  2. 2.
    DCC-NL: the portfolio in (21), where $\widehat{H}_{t}$ is estimated from the unconditional correlation matrix processed with nonlinear shrinkage, this procedure follows engle2019large.
  3. 3.
    DCS (tau): the portfolio in (21), where $\widehat{H}_{t}$ is estimated from the DCS model in Definition 3.2 using Kendall’s tau statistic.
  4. 4.
    DCS (rho): the portfolio in (21), where $\widehat{H}_{t}$ is estimated from the DCS model in Definition 3.2 using Spearman’s rho statistic.
  5. 5.
    $1/p$: the equal-weighted portfolio, a common benchmark advocated by demiguel2009optimal.
For the estimation of the inverse correlation matrix, we start from a rank-based SKEPTIC estimator constructed from Kendall’s tau and Spearman’s rho statistics. Rank-based dependence measures are substantially more robust than Pearson correlation under heavy tails, skewness, and other departures from Gaussianity, and they provide superior support recovery for the precision matrix in high-dimensional settings liu2012high. To obtain a stable and well-conditioned estimate of the inverse correlation matrix, we impose sparsity through the QUIC algorithm of hsieh2014quic, selecting the regularization parameter $\lambda$ via the StARS stability criterion of liu2010stability [^4] The tuning parameter is set to $\lambda=0.596$ for the S&P100, and $\lambda=0.547$ for the S&P500.. This approach ensures that only the most reliable conditional dependencies are retained and avoids applying shrinkage to a Pearson-based correlation matrix, which is known to be less reliable in non-Gaussian environments.
We estimate the portfolio weights using the five aforementioned covariance matrix estimators. We evaluate the portfolios using the following performance metrics. We include the annualized average value (AV), the annualized standard deviation (SD) of the out-of-sample portfolio returns, and the Sharpe ratio (SR), which is obtained by dividing the portfolio’s out-of-sample average value by its standard deviation (SR = AV/SD). We also include the turnover (TO) metric to capture the changes in portfolio weights, reflecting the magnitude of portfolio rebalancing costs:
$\text{TO}=\frac{1}{72}\sum_{h=2}^{73}|w_{h-1}-w_{h}|.$
S&P100 S&P500
AV SD SR TO AV SD SR TO
DCC 7.17 16.57 0.43 3.86 8.49 16.38 0.52 6.96
DCC-NL 7.49 16.42 0.46 3.57 8.96 15.70 0.57 5.10
DCS (tau) 10.00 16.45 0.61 0.47 7.68 15.81 0.48 0.83
DCS (rho) 10.95 17.05 0.64 0.37 9.23 17.13 0.54 0.52
EW 14.12 20.01 0.71 0.01 11.62 21.05 0.55 0.01
Table 3: Comparison of annualized performance measures (expressed in percentage) for S&P100 and S&P500 datasets. AV = Average Value, SD = Standard Deviation, SR = Sharpe ratio, TO = Turnover.
Table 3 reports the portfolio performance metrics. The key finding is the higher annualized return achieved by the DCS-based portfolios compared to the DCC-based ones, particularly for portfolios constructed from the S&P100 constituents. The slightly higher standard deviation of the DCS-based portfolios is compensated by a substantially higher return, resulting in a greater Sharpe ratio. Additionally, as is predictable, the portfolios constructed using the DCS model show a lower turnover. The ability of the SKEPTIC estimator to recover the true precision matrix in non-Gaussian cases can be exploited for portfolio construction to reduce transaction costs, without sacrificing the portfolio performance.
Nevertheless, for portfolios based on the S&P500 constituents, the Sharpe ratios of the DCS portfolios are slightly lower than those of the DCC-based and equally weighted portfolios. In contrast, while their standard deviations are not the lowest, they remain sufficiently contained to make the DCS-based portfolios a plausible and stable alternative, further supported by their lower turnover.
7.3 Portfolio Backtesting
We evaluate our portfolios in terms of market risk, as measured by the Value at Risk (VaR) and the Expected Shortfall (ES). Given a probability level $\alpha$, the VaR estimates the maximum potential daily loss at the specified confidence level $\alpha$, whereas the Expected Shortfall represents the expected loss conditional on the VaR threshold being exceeded.
Under the normality assumption, given the confidence level $\alpha$, the Value at Risk of the portfolio at $t+1$ is computed as:
$\operatorname{VaR}_{t+1}^{p}=-\sigma_{\operatorname{ptf},t+1}\cdot\Phi^{-1}(\alpha),$ (23)
and the related Expected Shortfall is given by:
$\mathrm{ES}=\sigma_{\operatorname{ptf},t+1}\frac{\varphi\left(\Phi^{-1}(\alpha)\right)}{\alpha},$ (24)
where $\sigma_{\operatorname{ptf},t+1}$ denotes the volatility of the portfolio estimated from the out-of-sample return analysis, and $\varphi$ the density of a standard normal distribution. To obtain the dynamic estimate of the portfolio volatility, we fit a GJR-GARCH(1,1) model to portfolio returns. We recall that the portfolio returns are defined as $r_{t+1}^{\mathrm{ptf}}=w_{t}^{\prime}r_{t+1}$, where $r_{t+1}$ are the realized returns of the stocks that compose the indices, and $r^{ptf}_{t+1}$ are the realized returns of the portfolios.
We set the confidence level at $\alpha=5\%$. To assess whether the model and estimation procedure produce reliable risk measure estimates, we conduct three backtesting procedures: the Unconditional Coverage (UC) test (kupiec1995techniques), the Conditional Coverage (CC) test (christoffersen2000relevant), and the Dynamic Quantile (DQ) test (engle2004caviar). These tests evaluate whether the empirical violation probability aligns with the expected violation probability implied by the confidence level, thereby capturing the realized risk. All three tests are based on the null hypothesis that the expected number of violations matches the observed number.
Table 4 reports the results of the backtesting procedures.
Table 4: Backtesting procedure results. In the Table, Fail no. indicates the percentage of failures, UC the test statistic values of the unconditional coverage test, $C$ the test statistic values of the conditional coverage test, and DQ the test statistic values of the dynamic quantile test. Note: $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
S&P 100 S&P 500
$VaR^{0.05}_{t+1}$ Fail no. UC $C$ DQ Fail no. UC $C$ DQ
DCC 0.05 0.26 0.38 7.54 0.05 0.30 0.35 3.14
DCC-NL 0.05 0.15 0.94 10.31 0.05 0.30 1.02 2.85
DCS Tau 0.05 0.01 0.02 5.14 0.05 0.26 0.73 3.88
DCS Rho 0.05 0.01 0.02 4.67 0.05 0.08 0.42 3.45
EW 0.05 0.54 0.60 6.25 0.05 0.26 0.95 5.30
$ES^{0.05}_{t+1}$ Fail no. UC $C$ DQ Fail no. UC $C$ DQ
DCC 0.03 10.19∗∗∗ 11.07∗∗∗ 15.40 0.03 24.99∗∗∗ 24.99∗∗∗ 23.88∗∗∗
DCC-NL 0.03 11.97∗∗∗ 12.09∗∗∗ 17.17∗∗ 0.03 19.59∗∗∗ 19.61∗∗∗ 20.69∗∗∗
DCS Tau 0.03 12.89∗∗∗ 13.09∗∗∗ 12.61 0.03 13.92∗∗∗ 14.13∗∗∗ 17.85∗∗
DCS Rho 0.03 13.88∗∗∗ 14.04∗∗∗ 13.53 0.03 11.06∗∗∗ 11.15∗∗∗ 13.50
EW 0.03 18.36∗∗∗ 20.84∗∗∗ 17.17∗∗ 0.03 19.59∗∗∗ 21.95∗∗∗ 20.83∗∗∗
The models deliver optimal VaR values, with the number of failures being equal to the expected 5% threshold. For the Expected Shortfall estimates, the values remain consistently below the 5% benchmark, thereby rejecting the null hypothesis of equality. The test statistic values, in conjunction with the observed number of violations, provide strong evidence of the reliability of our portfolios to effectively capture extreme tail events.
Figure 1: Backtesting results for SP100 and SP500 across the DCC, DCC-NL, DCS-Rho, DCS-Tau and EW specifications. Return observations are plotted against their corresponding VaR and ES estimates, with violations highlighted in green.
As illustrated in Figure 1, the DCS-Tau and DCS-Rho models exhibit a more flexible and responsive behaviour than both the DCC-based models and the EW benchmark, enabling them to capture shocks in portfolio returns promptly. This flexibility translates into a superior ability to model and predict systemic risk—measured through VaR and Expected Shortfall—with the DCS-Tau model delivering the best overall performance.
8 Conclusion
In this paper, we introduce the Dynamic Conditional SKEPTIC (DCS) model, an extension of the DCC model that relaxes the normality assumption by utilizing nonparametric distributions and rank-based statistics, such as Spearman’s rho and Kendall’s tau, to estimate the unknown correlation matrix.
We investigate the theoretical properties of the DCS model, including its stationarity and mixing properties, highlighting its robustness and flexibility in dynamically modeling the correlation matrix. The proposed model provides a more adaptable approach to capture complex market dynamics by addressing the limitations of the normality assumption in asset returns. A simulation study based on Montecarlo simulations shows the DCS in estimating the process parameters. The empirical analysis demonstrates that the DCS model achieves portfolio Sharpe ratios comparable to benchmark models while having lower turnover, thereby reducing portfolio costs. In the risk management application, the proposed estimators demonstrate optimal performance with a portfolio that, besides providing optimal average return and low rebalancing costs, ensures robust market risk tracking.
In summary, the Dynamic Conditional SKEPTIC model represents an appealing alternative to traditional methods for estimating the conditional correlation matrix, offering flexible assumptions and delivering interesting results for portfolios, particularly in the analysis of portfolios based on S&P100 constituents.
Appendix of Dynamic Conditional SKEPTIC
Appendix A Proof of Proposition 4.3
We recall the $\left\{W_{t}\right\}$ process in (10) that is a stationary Markov chain which implies $\textbf{X}_{t}$ a stationary process. It is known that its $\beta$-mixing coefficients satisfy:
$\beta_{k}=\frac{1}{2}\left\|F_{0,k}(a,b)-F(a)F(b)\right\|_{\mathrm{TV}},$ (25)
where $F_{0,k}$ is the joint distribution function of $X_{0}$ and $X_{k}$, and $\|\cdot\|_{\mathrm{TV}}$ is total variation, in the Vitali sense. We revisit the result of beare2010copulas. From Sklar’s theorem, and liu2009nonparanormal we have
$\beta_{k}$ $=\frac{1}{2}\|F_{0,k}(a,b)-F(a)F(b)\|_{TV}$
$=\frac{1}{2}\|\Phi_{\rho_{k}}(\Phi^{-1}(a),\Phi^{-1}(b))-\Phi(a)\Phi(b)\|_{TV}$
$\leq\frac{1}{2}\|\Phi_{\rho_{k}}(\Phi^{-1}(a),\Phi^{-1}(b))-ab\|_{TV},$
where $\rho_{k}$ is the correlation between $X_{0}$ and $X_{k}$ that depends only to $k$. Let $\varphi_{k}$ denote the density of the semiparametric Gaussian distribution as defined in (7). This function is differentiable and continuous. It holds that $\beta_{k}\leq\frac{1}{2}\left\|\varphi_{k}-1\right\|_{1},$ and consequently, $\beta_{k}\leq\frac{1}{2}\left\|\varphi_{k}-1\right\|_{2}.$ Since $\varphi_{k}$ is a symmetric, square-integrable joint density function with uniform marginals, it admits a mean-square convergent expansion:
$\varphi_{k}(a,b)=1+\sum_{i=1}^{\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(b),$
where the eigenvalues $\left\{\lambda_{i}\right\}$ form a nonincreasing square-summable sequence of non-negative real numbers and the eigenfunctions $\left\{\phi_{i}\right\}$ form a complete orthonormal sequence in $L_{2}[0,1]$. We know from beare2010copulas that:
$C_{k+1}(a,b)=\int_{0}^{1}\frac{\partial C_{k}(a,v)}{\partial v}\cdot\frac{\partial C_{1}(v,b)}{\partial v}\,dv,$ (26)
which allows us to infer the spectral expansion of the density $\varphi_{k}$:
$\varphi_{k}(a,b)=1+\sum_{i=1}^{\infty}\lambda_{i}^{k}\phi_{i}(a)\phi_{i}(b).$ (27)
We now have
$\left\|c_{k}-1\right\|_{2}=\left\|\sum_{i=1}^{\infty}\lambda_{i}^{k}\phi_{i}(a)\phi_{i}(b)\right\|_{2},$ (28)
so with two applications of Parseval’s equality, we obtain
$\left\|c_{k}-1\right\|_{2}=\left(\sum_{i=1}^{\infty}\lambda_{i}^{2k}\right)^{\frac{1}{2}}\leq\lambda_{1}^{k-1}\left(\sum_{i=1}^{\infty}\lambda_{i}^{2}\right)^{\frac{1}{2}}=\lambda_{1}^{k-1}\|c-1\|_{2},$ (29)
where $\lambda_{1}$ is equal to the maximal correlation of $C$. Since this quantity is assumed to be less than 1, the proof is complete.
Appendix B Proof of Proposition 4.5
Let $\widehat{R}^{\tau}_{t}$ and $\widehat{R}^{\rho}_{t}$ denote the estimated correlation matrices based on Kendall’s tau and Spearman’s rho, respectively. For notational simplicity, we omit the time dependence on $t$ without loss of generality. By Lemma 3.1, we have:
$R=\sin\left(\frac{\pi}{2}R^{\tau}\right)=2\sin\left(\frac{\pi}{6}R^{\rho}\right).$
We exploit the Lipschitz continuity of the estimators and align our proof to the results of han2018exponential. We recall that let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, respectively. Denote the Gaussian cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$ by $\Phi_{i}$ and $\Phi_{j}$. The population version of Spearman’s rho is: $R^{\rho}]_{ij}=\operatorname{Corr}\left(\Phi_{i}\left(X_{i}\right),\Phi_{j}\left(X_{j}\right)\right)$, and the population version of Kendall’s tau is:
$R^{\tau}]_{ij}=\operatorname{Corr}\left(\operatorname{sign}\left(X_{i}-\widetilde{X}_{i}\right),\operatorname{sign}\left(X_{j}-\widetilde{X}_{j}\right)\right)$. We establish a bound for Spearman’s rho:
$\mathbb{P}\left(\left\|2\sin\left(\frac{\pi}{6}\widehat{R}^{\rho}\right)-R\right\|_{\max}>x+C/\sqrt{T}\right)=$ (30)
$=\mathbb{P}\left(\left\|2\sin\left(\frac{\pi}{6}\widehat{R}^{\rho}\right)-2\sin\left(\frac{\pi}{6}{R}^{\rho}\right)\right\|_{\max}>x+C/\sqrt{T}\right)$
$\leq\frac{p(p-1)}{2}\max_{i,j}\mathbb{P}\left(\left|\widehat{R}_{ij}^{\rho}-R^{\rho}_{ij}\right|>x+C/\sqrt{T}\right)$
Notice that:
$\left|3\tau_{2}\left(\left(\boldsymbol{X}_{1}\right)_{\{i,j\}},\ldots,\left(\boldsymbol{X}_{T}\right)_{\{i,j\}}\right)/(T+1)\right|\leq\frac{3}{T}$
and
$\rho_{3}\left(\left(\boldsymbol{X}_{1}\right)_{\{i,j\}},\ldots,\left(\boldsymbol{X}_{T}\right)_{\{i,j\}}\right)$
has a kernel upper bound by an absolute constant $M_{0}$. By choosing $C$ in (30) large enough, then
$\mathbb{P}\left(\|\widehat{R}^{\rho}-R^{\rho}\|_{\max}>x+C/\sqrt{T}\right)\leq p^{2}\exp\left(-\frac{C^{\prime}x^{2}T}{M_{0}^{2}+M_{0}x(\log T)(\log\log 4T)}\right).$
Picking $x=\left(3M_{0}^{2}\log(Tp)/\left(C^{\prime}T\right)\right)^{1/2}$, we have
$\mathbb{P}\left(\|\widehat{R}^{\rho}-R^{\rho}\|_{\max}>x+C/\sqrt{T}\right)\rightarrow 0.$
For Kendall’s tau, we have:
$\mathbb{P}\left(\|\sin\left(\frac{\pi}{2}\widehat{R}^{\tau}\right)-R\|_{\max}>\frac{C}{\sqrt{T}}\right)$
$\mathbb{P}\left(\|\sin\left(\frac{\pi}{2}\widehat{R}^{\tau}\right)-\sin\left(\frac{\pi}{2}R^{\tau}\right)\|_{\max}>\frac{C}{\sqrt{T}}\right)$
$\leq\frac{p(p-1)}{2}\mathbb{P}\left(\max_{ij}\left|\widehat{R}^{\tau}_{ij}-R^{\tau}_{ij}\right|>\frac{C}{\sqrt{T}}\right)$
$\leq p^{2}\exp\left(-\frac{C}{\sqrt{T}}\right).$
Appendix C Simulation Results
Tables 5, 6, and 7 summarize the results of the simulations. The tables consider respectively 1%, 5% and 10% of data contamination.
Table 5: Mean and standard deviation of the estimated parameter $\alpha$ and $\beta$ assuming $\delta=0.01$.
$\delta=0.01$ DCC DCC NL DCS Rho DCS Tau
$T$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$
$p=50$
500 0.028 0.928 0.028 0.928 0.035 0.926 0.035 0.924
s.d. 0.004 0.013 0.004 0.013 0.005 0.006 0.005 0.006
1000 0.025 0.941 0.026 0.942 0.029 0.939 0.029 0.938
s.d. 0.001 0.003 0.001 0.003 0.001 0.001 0.001 0.002
2000 0.025 0.943 0.025 0.943 0.028 0.942 0.028 0.940
s.d. 0.000 0.002 0.000 0.002 0.002 0.005 0.002 0.005
$p=97$
500 0.028 0.928 0.028 0.928 0.035 0.926 0.035 0.924
s.d. 0.004 0.013 0.004 0.013 0.005 0.006 0.005 0.006
1000 0.025 0.941 0.026 0.942 0.029 0.939 0.029 0.938
s.d. 0.001 0.003 0.001 0.003 0.001 0.001 0.001 0.002
2000 0.025 0.943 0.025 0.943 0.028 0.942 0.028 0.940
s.d. 0.000 0.002 0.000 0.002 0.002 0.005 0.002 0.005
$p=250$
500 0.015 0.945 0.015 0.945 0.023 0.930 0.023 0.925
s.d. 0.004 0.010 0.004 0.010 0.002 0.008 0.002 0.008
1000 0.016 0.947 0.016 0.947 0.022 0.941 0.022 0.939
s.d. 0.001 0.002 0.001 0.002 0.001 0.003 0.001 0.003
2000 0.016 0.951 0.016 0.951 0.020 0.949 0.020 0.948
s.d. 0.001 0.005 0.001 0.005 0.001 0.001 0.001 0.001
$p=497$
500 0.009 0.955 0.010 0.955 0.019 0.942 0.019 0.938
s.d. 0.002 0.003 0.002 0.003 0.002 0.001 0.002 0.001
1000 0.014 0.944 0.014 0.944 0.018 0.945 0.018 0.943
s.d. 0.002 0.003 0.002 0.002 0.001 0.006 0.001 0.006
2000 0.013 0.955 0.013 0.955 0.018 0.951 0.018 0.950
s.d. 0.001 0.002 0.001 0.002 0.001 0.001 0.001 0.001
Table 6: Mean and standard deviation of the estimated parameter $\alpha$ and $\beta$ assuming $\delta=0.05$.
$\delta=0.05$ DCC DCC NL DCS Rho DCS Tau
$T$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$
$p=50$
500 0.024 0.936 0.024 0.937 0.031 0.929 0.031 0.926
s.d. 0.002 0.007 0.002 0.007 0.001 0.007 0.001 0.008
1000 0.024 0.934 0.024 0.934 0.027 0.936 0.027 0.933
s.d. 0.004 0.006 0.004 0.006 0.003 0.010 0.003 0.010
2000 0.025 0.943 0.025 0.943 0.028 0.943 0.028 0.942
s.d. 0.001 0.001 0.001 0.001 0.001 0.001 0.001 0.001
$p=97$
500 0.021 0.930 0.022 0.930 0.026 0.935 0.027 0.930
s.d. 0.003 0.008 0.003 0.008 0.003 0.008 0.003 0.009
1000 0.023 0.935 0.023 0.935 0.025 0.937 0.025 0.935
s.d. 0.001 0.002 0.001 0.002 0.001 0.003 0.001 0.003
2000 0.019 0.947 0.020 0.947 0.023 0.947 0.023 0.946
s.d. 0.001 0.003 0.001 0.003 0.001 0.005 0.001 0.005
$p=250$
500 0.014 0.941 0.014 0.942 0.020 0.937 0.020 0.932
s.d. 0.005 0.018 0.005 0.018 0.002 0.002 0.002 0.002
1000 0.014 0.948 0.014 0.948 0.020 0.942 0.020 0.940
s.d. 0.002 0.008 0.002 0.007 0.001 0.003 0.001 0.003
2000 0.012 0.957 0.012 0.957 0.020 0.951 0.020 0.950
s.d. 0.005 0.009 0.005 0.009 0.002 0.004 0.002 0.004
$p=497$
500 0.009 0.950 0.010 0.950 0.020 0.931 0.020 0.926
s.d. 0.004 0.010 0.004 0.010 0.001 0.005 0.001 0.004
1000 0.013 0.948 0.013 0.949 0.019 0.946 0.019 0.944
s.d. 0.001 0.002 0.001 0.002 0.001 0.001 0.001 0.002
2000 0.013 0.953 0.013 0.954 0.017 0.951 0.017 0.950
s.d. 0.001 0.001 0.001 0.002 0.001 0.001 0.001 0.001
Table 7: Mean and standard deviation of the estimated parameter $\alpha$ and $\beta$ assuming $\delta=0.1$.
$\delta=0.1$ DCC DCC NL DCS Rho DCS Tau
$T$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$
$p=50$
500 0.025 0.930 0.026 0.931 0.032 0.923 0.032 0.917
s.d. 0.002 0.005 0.002 0.005 0.003 0.008 0.003 0.007
1000 0.026 0.933 0.026 0.933 0.031 0.929 0.031 0.928
s.d. 0.002 0.010 0.002 0.010 0.002 0.014 0.002 0.014
2000 0.022 0.946 0.022 0.946 0.027 0.942 0.027 0.941
s.d. 0.003 0.007 0.003 0.007 0.002 0.003 0.002 0.002
$p=97$
500 0.021 0.931 0.021 0.931 0.029 0.923 0.029 0.919
s.d. 0.003 0.012 0.003 0.012 0.004 0.007 0.005 0.010
1000 0.020 0.947 0.020 0.947 0.023 0.946 0.023 0.944
s.d. 0.001 0.003 0.001 0.003 0.001 0.002 0.001 0.002
2000 0.019 0.948 0.019 0.948 0.023 0.947 0.023 0.946
s.d. 0.002 0.004 0.002 0.004 0.001 0.003 0.001 0.003
$p=250$
500 0.012 0.947 0.012 0.947 0.021 0.933 0.021 0.929
s.d. 0.001 0.003 0.001 0.003 0.002 0.008 0.003 0.010
1000 0.014 0.945 0.014 0.946 0.020 0.947 0.020 0.944
s.d. 0.003 0.014 0.003 0.014 0.003 0.004 0.003 0.004
2000 0.015 0.951 0.015 0.951 0.018 0.951 0.018 0.950
s.d. 0.001 0.003 0.001 0.003 0.000 0.001 0.000 0.001
$p=497$
500 0.012 0.947 0.012 0.947 0.021 0.933 0.021 0.929
s.d. 0.001 0.003 0.001 0.003 0.002 0.008 0.003 0.010
1000 0.014 0.945 0.014 0.946 0.020 0.947 0.020 0.944
s.d. 0.003 0.014 0.003 0.014 0.003 0.004 0.003 0.004
2000 0.015 0.951 0.015 0.951 0.018 0.951 0.018 0.950
s.d. 0.001 0.003 0.001 0.003 0.000 0.001 0.000 0.001
Appendix D Tests of Portfolio Performance Metrics
In this section, we perform statistical tests to assess the significance of the portfolio metrics.
D.1 t-test for Average Sample Means
Let $\pi_{r}$ and $\pi_{s}$ represent two generic portfolios, both with $n$ observations. We perform the two-sample $t$-test. Let $r_{t,\pi_{r}}$ and $r_{t,\pi_{s}}$ denote the log-return at time $t$ of portfolio $\pi_{r}$ and $\pi_{s}$. The two-sample $t$-test is defined as:
$\hat{t}_{\pi_{r},\pi_{s}}=\frac{\bar{r}_{\pi_{r}}-\bar{r}_{\pi_{s}}}{\sqrt{\frac{s_{\pi_{r}}^{2}+s_{\pi_{s}}^{2}}{n}}}$
where $\bar{r}_{\pi_{r}}$ and $\bar{r}_{\pi_{s}}$ are, respectively, the average returns of $\pi_{r}$ and $\pi_{s}$. The values $s_{\pi_{r}}^{2}$ and $s_{\pi_{s}}^{2}$ are the sample variance of the portfolios, defined as:
$s_{\pi,r}^{2}=\frac{1}{n-1}\sum_{t-1}^{n}\left(r_{\pi_{r,t}}-\bar{r}_{\pi_{r}}\right)^{2},\quad s_{\pi,s}^{2}=\frac{1}{n-1}\sum_{t-1}^{n}\left(r_{\pi_{s,t}}-\bar{r}_{\pi_{s}}\right)^{2}.$ (31)
In Table 8, we present the results of the $t$ test.
SP100 SP500
Comparison $t$-Test Value $t$-Test Value
DCC vs DCS (tau) -0.354 0.087
DCC vs DCS (rho) -0.396 -0.075
DCC vs DCC NL -0.034 -0.052
DCC vs EW -0.662 -0.279
DCC NL vs EW -0.634 -0.239
DCC NL vs DCS (rho) -0.364 -0.026
DCC NL vs DCS (tau) -0.321 0.142
DCS (tau) vs DCS (rho) -0.043 -0.162
DCS (tau) vs EW -0.335 -0.359
DCS (rho) vs EW -0.295 -0.209
Table 8: Summary of the $t$-tests of annualized average sample means. In the Table, $t$-statistic values are shown. Note: $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
D.2 F-test for Sample Variance
We recall from (31) $s_{\pi_{r}}^{2}$, which is the unbiased estimator of variance, and we perform the $F$-test on the ratio of variances:
$F=\frac{s_{\pi_{r}}^{2}}{s_{\pi_{s}}^{2}},\quad H_{0}:F=1$
In Table 9, the results of the $F$-test for the variances of the portfolios are shown.
SP100 SP500
Comparison $F$-Test Value $F$-Test Value
DCC vs DCS (tau) 0.947 1.068
DCC vs DCS (rho) 0.942 0.924
DCC vs DCC NL 1.018 $1.089^{*}$
DCC vs EW $0.686^{***}$ $0.62^{***}$
DCC NL vs EW $0.674^{***}$ $0.57^{***}$
DCC NL vs DCS (rho) 0.926 $0.849^{***}$
DCC NL vs DCS (tau) 0.93 0.981
DCS (tau) vs DCS (rho) 0.995 $0.865^{***}$
DCS (tau) vs EW $0.724^{***}$ $0.581^{***}$
DCS (rho) vs EW $0.728^{***}$ $0.671^{***}$
Table 9: Summary of the $F$-tests of annualized rate of variances. In the Table, $F$-statistic values are shown. Note: $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
D.3 Sharpe ratio Test
We perform the Sharpe ratio test in the ”PeerPerformance” R package described in ardia2018peer using the block bootstrap approach. Table 10 presents the results of the Sharpe ratio test.
SP100 SP500
Comparison $\Delta$ Sharpe ratio $\Delta$ Sharpe ratio
DCC vs DCS (tau) 0.012-0.012 $0.002$
DCC vs DCS (rho) 0.013-0.013 0.001-0.001
DCC vs DCC NL $0.001$ $0.003$
DCC vs EW 0.017-0.017 0.002-0.002
DCC NL vs EW 0.016-0.016 $0.001$
DCC NL vs DCS (rho) 0.012-0.012 $0.002$
DCC NL vs DCS (tau) 0.010-0.010 $0.005$
DCS (tau) vs DCS (rho) $0.001$ $0.004$
DCS (tau) vs EW 0.006-0.006 0.004-0.004
DCS (rho) vs EW 0.004-0.004 0.001-0.001
Table 10: Summary of the ardia2018peer test for Sharpe ratio. The table reports differences in the annualized Sharpe ratio values. Note: $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.